OpenStudy (chaise):

Can anyone help me integrate this manually? sqrt(h)/(0.1-sqrt(h)) I have a feeling you have to use polynomial long division, but I don't get the same answer as a calculator. If anyone can shed any light that would be awesome.

OpenStudy (chaise):

www.wolframalpha.com/input/?i=integal+of+sqrt(h)%2F(0.1-sqrt(h))

OpenStudy (yttrium):

You can substitute \[u=\sqrt{h}\]

OpenStudy (chaise):

Please continue, I have done this already

OpenStudy (yttrium):

Show us what youv'e got. And let us verify.

OpenStudy (chaise):

Can you pls just do it for me?

OpenStudy (yttrium):

What is du, then?

OpenStudy (chaise):

dh=du2sqrt(h) or du/dh=1/2sqrt(h)

OpenStudy (chaise):

Please just do this question for me, haha

OpenStudy (chaise):

I will understand when I see the process.

OpenStudy (yttrium):

Okay. So\[dh = 2\sqrt{h} du\] Upon substitution, we'll get \[2\int\limits_{}^{}\frac{ u^2 }{ 0.1-u }du\] Then perform the long division. I think you can do this right?

OpenStudy (chaise):

Okay, I think after long division I should get.... Can you confirm? integral -u-0.1+(0.01/-u+0.1).du

OpenStudy (yttrium):

Correct! But you can actually rewrite this like 2\[2\int\limits_{}^{}(-u- \frac{0.01 }{ u-0.1 }-0.1) du\] Just integrate them separately. :D

OpenStudy (chaise):

Thank you so much. Awesome.

OpenStudy (yttrium):

Easy, right?! haha. :)

OpenStudy (chaise):

Uhm, not really. I'm just going with it, haha.

OpenStudy (yttrium):

Just continue where you are. :)

OpenStudy (chaise):

Oh, I can integrate the polynomial, haha. Long division and the substitutiuon was hard

OpenStudy (yttrium):

Yeah. As what I've told you. :)